You wish to make a 0.380 M hydrobromic acid solution from a stock solution of 3.00 M hydrobromic acid. How much concentrated acid must you add to obtain a total volume of 75.0 mL of the dilute solution?

________ M.
1
Expert's answer
2019-03-18T02:19:20-0400

"C_m= \frac{m}{M \cdot V}"

Cm - molarity or molar concentration of the solution

m - mass of the solute

M - molar mass of the solute

V - volume of the solution


"\frac{m(HBr)}{M(HBr)}=const=C_{m1} \cdot V_1=C_{m2} \cdot V_2"

Cm1 - initial molarity of the solution

V1 - initial volume of the solution

Cm2 - final molarity of the solution

V2 - final volume of the solution

"V_1=\frac{C_{m2} \cdot V_2}{C_{m1}}"

"V_1=\frac{0.380 \; M \cdot 75.0 \; mL}{3.00 \; M}=9.5 \; mL"


Answer: 9.5 mL.